# Signatures of cubic gravity in the strong regime

### v1

@@ II.6 Regimes of validity
Figure 2: Solutions with a positive coupling constant. Numerical solutions (black lines) are shown for constants $C_{0}$ fixed to the mass of SgrA and $\Lambda_{\text{Eff}}=\Lambda_{\text{CDM}}$, where the value of $\Lambda$ is determined using Eq. (16). The solutions correspond to two values of the coupling constant: $\lambda=0.05[M_{\text{Pl}}^{5}r_{s}^{4}]$ (left panel) and $\lambda=0.1[M_{\text{Pl}}^{5}r_{s}^{4}]$ (right panel). The blue and red solid lines represent the asymptotic solution Eq. (15) and the weak coupling solution Eq. (12), respectively. As expected, the former agrees well for large values of $r$, while the latter is applicable for small values of the coupling constant $\lambda$. The inset in the left panel compares our numerical solution with the Kottler and Schwarzschild metrics, revealing a smaller event horizon than those predicted by these Einstein equation solutions. The right panel highlights the existence of solutions with a naked singularity, as demonstrated by the inset, which illustrates the divergence of the Ricci scalar at $r=0$.
@@
In previous sections, we presented and discussed the approximated solutions (12), (15), and (20). Here, we analyze the validity of these solutions by comparing them with the full numerical solutions. To facilitate this comparison, we introduce the following dimensionless variables:
@@
into Eq. (II). Notice that for $C_{0}=1/2$ and $M_{\text{Pl}}=1$ the new and old variables coincide. For simplicity, we omit the bars in the notation throughout the rest of the paper and when dimensionful quantities are required, they will be specified explicitly.
@@
Figure 2 presents the numerical solution for two possible values of the positive coupling constant: $\lambda=0.05$ (left panel) and $\lambda=0.1$ (right panel). We used an adaptive explicit fifth-order Runge-Kutta routine for the numerical integration, with errors estimated using a fourth-order routine [36, 37, 38]. To ensure that the solutions are ghost-free (see section II.3 for a discussion) and reproduce the expected asymptotic behavior, the value of $\Lambda_{\text{Eff}}$ is fixed to the observational cosmological constant $\Lambda_{\text{CDM}}$ value. The constant $C_{0}$ is taken as the mass of SgrA, $C_{0}=4\times 10^{6}\,M_{\odot}$ [39]. As shown, the approximate series solutions (12) (up to $\mathcal{O}(\epsilon^{4})$) and (15) (up to $\mathcal{O}(\lambda^{2},C_{0}^{3}/r^{9})$) agree with the numerical results for large values of $r$. In both cases, the validity of the series deteriorates as they approach the horizon, a behavior consistent with the assumptions used in their construction. However, it can also be observed that, as expected, the asymptotic series exhibits a wider range of validity compared to the weak coupling as the $\lambda$ value increases. On the other hand, for comparison, the corresponding Schwarzschild solution (with $M=C_{0}$) and Kottler solution (with $M=C_{0},\Lambda=\Lambda_{\text{CDM}}$) are shown in the inset of the left panel. As noted and predicted in the previous section (for $\lambda>0$), the non-cosmological horizon is smaller than those of the Schwarzschild and Kottler solutions. In the right panel of Fig. 2, an illustrative solution with only one horizon (the cosmological horizon) is shown. Although the profile is regular at the origin, this solution contains a naked singularity, as the Ricci scalar diverges at $r=0$ (see the inset). Our numerical results indicate the existence of this type of solution when $\lambda\gtrapprox 0.09$ (with $C_{0},\Lambda_{\text{Eff}}$ fixed to the values specified in the figure). Finally, the approximate near horizon solution Eq. (20) is not shown because it is not valid in this branch (see Fig. 1).
Figure 3 presents a similar analysis (again fixing $\Lambda_{\text{eff}}$ to the cosmological constant value), but this time for a negative coupling constant $\lambda=-5$. The left panel shows an illustrative (non-physical) solution to demonstrate the methodology used: connecting the near and asymptotic regions by the numerical solution (black line). The region near the horizon is described by the solution Eq. (20) (blue line), while for large $r$, the solution matches the asymptotic approximation Eq. (15) (red line). Our results indicate (and validate our analytical findings) that, for this branch, the exterior horizons are larger than those of the respective Schwarzschild and Kottler solutions (see the corresponding insets). The right panel shows the results obtained by fixing the constant $C_{0}$ as the mass of SgrA.
@@
III Possibles signatures of the cubic contribution in light deflection
Figure 4: Schematic and illustrative null geodesic triangle array. The left panel displays a schematic representation of a triangular array with vertices $\phi_{p}$, bounded by three null geodesics $\mathcal{C}_{p}$ ($p=1,2,3$) localized in the equatorial plane of a Riemannian space $\mathcal{M}_{\text{opt}}$ defined by the optical metric. Each geodesic $\mathcal{C}_{p}$ is associated with an impact parameter $b_{p}$ at an angular position $\phi_{b_{p}}$. For all triangular arrays explored in this paper, the vertices are fixed at: $\phi_{1}=\pi-\phi_{\text{Min}},\phi_{2}=\phi_{\text{Min}}$, and $\phi_{3}=\pi/2$, where $\phi_{\text{Min}}=0.35$rad. The corresponding internal angles $\beta_{p}$, are used to compute the angular difference Eq. (28). The central and right panels display triangle arrays where the spacetime under study is given by the numerical solutions of Eq. (II) with a positive coupling constant $\lambda=0.05$ (blue dashed line), and the background is described by the Kottler solution (solid black line). The central panel illustrates how the triangle array becomes more deformed due to its proximity to the gravitational source. In contrast, the right panel shows the same triangular array positioned much farther from the gravitational source. The insets reveal that, although these triangles appear similar, they exhibit minor discrepancies attributable to the $\lambda$ effect. Further details and discussions are provided in the main text.
In this section, we examine the angular difference $\alpha$, as proposed in Ref. [21], within the framework of Einstein cubic gravity. As mentioned in the introduction, this angular difference provides an alternative method for determining the deflection angle in static, spherically symmetric, and non-asymptotically flat spacetimes [40]. The method involves calculating the internal angles of a closed polygon (specifically, a triangle) formed by null geodesics in the spacetime of interest, which encode the effects of local curvature, and comparing them to a reference polygon (triangle) constructed in a background spacetime. A schematic representation of the construction of this triangle is shown in the left panel of Fig. 4.
As derived in Ref. [21], the angle $\alpha$ is determined from the angular difference between the internal angles of the triangular array:
@@ III Possibles signatures of the cubic contribution in light deflection
where $\beta_{p}^{(i)}$ are the internal angles of the triangle placed in the spacetimes: $i=1$ the spacetime of interest (associated with the modified gravity) and $i=2$ the spacetime that serves as background. For instance, if $i=1$ is the Schwarzschild spacetime and $i=2$ is Minkowski, then in the limit where the impact parameters of the null geodesics $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ (see Fig. 4) become sufficiently large (i.e., $b_{2}=b_{3}\to\infty$), the textbook formula for the deflection angle of the Schwarzschild metric is recovered.666 To review the procedure for obtaining the textbook formula for the deflection angle of the Schwarzschild metric, we refer the reader to Ref. [41], Section $7.1$ of Ref. [42], or Section $5$ of Ref. [43]. That is, the triangular configuration behaves as if the light source and the observer were located in the flat region at spatial infinity [21].
@@ III.1 Triangular numerical construction
where $b:=L/E=[r^{2}/f(r)]d\phi/dt$ is the impact parameter, defined in terms of the conserved photon energy $E=f(r)dt/d\lambda$ and angular momentum $L=r^{2}d\phi/d\lambda$, with $\lambda$ an affine parameter along the light trajectory. In terms of the dimensionless variables (27) and introducing the substitution $u:=1/r$, Eq. (29) can be rewritten as:
@@
The second step consists in solving Eq. (30) with appropriate boundary conditions for the three null geodesics $C_{1},C_{2}$, and $C_{3}$, such that a triangular configuration is formed. For simplicity, the vertices are chosen as:
@@
as shown in the left panel of Fig. 4. For the geodesic $\mathcal{C}_{1}$, we apply a Neumann boundary condition, $du/d\phi=0$ at $\phi=\pi/2$. For the remaining geodesics, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$, we use Dirichlet boundary conditions such that the bottom vertex of $\mathcal{C}_{2}$ ($\mathcal{C}_{3}$) coincides with the right (left) vertex of $\mathcal{C}_{1}$ at $\phi_{2}=\phi_{\text{Min}}$ ($\phi_{3}=\phi_{\text{Max}}$), where $[\phi_{\text{Min}},\phi_{\text{Max}}]$ is the angular domain of integration for $\mathcal{C}_{1}$. For simplicity, the impact parameters $b_{2}$ and $b_{3}$ of geodesics $\mathcal{C}_{2}$ and $\mathcal{C}_{2}$ respectively, are taken to be equal and fixed to $1.5b_{1}$. We only vary the parameter $b_{1}$, which corresponds to the geodesic $\mathcal{C}_{1}$ that describes the closest trajectory to the gravitational source. The numerical integration is carried out using the same methodology as in Sec. II.6.
@@
Once the triangles are constructed, we compute their internal angles $\beta_{p}$ and, from these, the angular difference $\alpha$ as the final step. The internal angles are obtained from [21]
@@
III.2 Cubic terms contributions
@@
$b_{1}[r_{s}]$ $-5.0$ $-1.5$ $0.05$ $0.1$
$10$ $6.46\times 10^{3}$ $5.65\times 10^{3}$ $0.61\times 10^{2}$ $1.22\times 10^{2}$
@@
As initial examples, we considered as the spacetime of interest the (numerical and approximated) solutions of Einstein cubic gravity with the integration constant $C_{0}=4.3\times 10^{6}M_{\odot}$ (corresponding to the observed mass of SgrA) and coupling constant values: $\lambda=-5.0,-1.5,0.05,0.1$. The impact parameters of the triangular array were chosen such that the relation $b_{2}=b_{3}=1.5b_{1}$ holds, while $\phi_{\text{min}}$ was fixed at $0.3$ rad. Table 2 presents the angular difference $\alpha$ computed using the Kottler (and Schwarzschild) solution as the background. As noted above, in this scenario, the values of $\alpha$ encode the (gravitational) effect of the cubic terms due to the mass contribution. Our results indicate that as $|\lambda|\to 0$ (for a fixed $b_{1}$), the value of $\alpha$ either diminishes or remains constant. This behavior highlights that the gravitational contribution of the cubic terms due to a massive source becomes negligible as $|\lambda|$ decreases. In the limit $\lambda=0$, it is expected that $\alpha=0$. Consequently, the Kottler/Schwarzschild solution is recovered, corresponding to the regime where general relativity is restored. The cases where $\alpha$ remains constant despite the value of $\lambda$ correspond to triangular arrays that are sufficiently far from the gravitational source (e.g. $b_{1}=10^{4}\,r_{s}$), such that the correction introduced by the coupling constant term in the asymptotic solution, Eq. (15), becomes negligible. In contrast, the value of the angular difference increases considerably when the coupling constant $\lambda$ is fixed and the impact parameter $b_{1}$ decreases. This is a direct consequence of the triangular configurations being closer to the gravitational source, which makes the contribution of the cubic terms more relevant.
@@
As a second example, we fixed the impact parameters $b_{1}=4r_{s}$, $b_{2}=b_{3}=1.5b_{1}$, and varied the coupling constant $\lambda$ within the interval $[0,10]$ taking $C_{0}=4.3\times 10^{6}M_{\odot}$. The results are shown in Fig. 5. As observed from the left panel, when $\lambda\gtrapprox 0.1$, the solution does not present an black hole horizon (although the cosmological horizon remains present). This indicates that the value of $\lambda$ is constrained to $\lambda\lessapprox 0.1$ in order to ensure the presence of an exterior event horizon in black holes like SgrA (keeping in mind that the mass that we are using is very close to that of SgrA). The right panel shows the angular difference for these solutions. As observed, this increases as $\lambda$ grows. However, the results seem to indicate that it asymptotically approaches a certain value. We do not delve deeper into identifying it, as it corresponds to non-physically viable values of $\lambda$ due to the large modifications to the spacetime metric near the massive source, leading even to the disappearance of the event horizon.
@@
A similar analysis to the previous one is presented in Fig. 6, but in this case for $\lambda\in[-10,0]$ and $b_{1}=10r_{s}$. The last change is due to the exterior horizon moving to the right for this branch, and we need to ensure that the triangular array is not inside it. In this case, we always have an exterior event horizon (greater than $2C_{0}/M_{\textrm{Pl}}^{2}$), which could, in principle, reproduce black hole types like SgrA. As observed from the left panel, when $\lambda$ decreases (increases its magnitude $|\lambda|$), the horizon’s radius shifts to the right. However, this shift becomes progressively smaller; for example, the horizon displacement from a solution with $\lambda=-1$ to a solution with $\lambda=-2$ is approximately $0.2r_{s}$, while for $\lambda=-9$ to $\lambda=-10$ it is approximately $0.04r_{s}$. The right panel shows the angular difference. In this case, our results seem to indicate a linear increase (the $y$-axis is on a logarithmic scale). This is a direct consequence of the horizon approaching to the position of the triangular array. As discussed in the first example, this implies that the effect of the cubic terms becomes more relevant, and the differences relative to the triangular background array become more significant.

### v2

@@ II.6 Regimes of validity
Figure 2: Solutions with a positive coupling constant. Numerical solutions (black lines) are shown for constants $C_{0}$ fixed to the mass of SgrA and $\Lambda_{\text{Eff}}=\Lambda_{\text{CDM}}$, where the value of $\Lambda$ is determined using Eq. (16). The solutions correspond to two values of the coupling constant: $\lambda=0.05[M_{\text{Pl}}^{5}r_{s}^{4}]$ (left panel) and $\lambda=0.1[M_{\text{Pl}}^{5}r_{s}^{4}]$ (right panel). The blue and red solid lines represent the asymptotic solution Eq. (15) and the weak coupling solution Eq. (12), respectively. As expected, the former agrees well for large values of $r$, while the latter is applicable for small values of the coupling constant $\lambda$. The inset in the left panel compares our numerical solution with the Kottler and Schwarzschild metrics, revealing a smaller event horizon than those predicted by these solutions. The right panel highlights the existence of solutions with a naked singularity, as demonstrated by the inset, which illustrates the divergence of the Ricci scalar at $r=0$.
@@
In previous sections, we presented and discussed the approximated solutions (12), (15), and (20). Here, we analyze the validity of these solutions by comparing them with the full numerical solutions. To facilitate this comparison, we introduce the following dimensionless variables,
@@
into Eq. (II). Notice that for $C_{0}=1/2$ and $M_{\text{Pl}}=1$ the new and old variables coincide. For simplicity, we omit the bars in the notation throughout the rest of the paper and when dimensionful quantities are required, they are specified explicitly.
@@
Figure 2 presents the numerical solutions for two possible values of the positive coupling constant: $\lambda=0.05$ (left panel) and $\lambda=0.1$ (right panel). We used an adaptive explicit fifth-order Runge-Kutta routine for the numerical integration, with errors estimated using a fourth-order routine [37, 38, 39]. To ensure that the solutions are ghost-free (see Section II.3 for a discussion) and reproduce the expected asymptotic behavior, the value of $\Lambda_{\text{Eff}}$ is fixed to the observational cosmological constant $\Lambda_{\text{CDM}}$. The constant $C_{0}$ is taken as the mass of SgrA, $C_{0}=4\times 10^{6}\,M_{\odot}$ [40]. As shown, the approximate series solutions (12) (up to $\mathcal{O}(\epsilon^{4})$) and (15) (up to $\mathcal{O}(\lambda^{2},C_{0}^{3}/r^{9})$) agree with the numerical results for large values of $r$. In both cases, the validity of the series deteriorates as they approach the horizon, a behavior consistent with the assumptions used in their construction. However, it can also be observed that, as expected, the asymptotic series exhibits a wider range of validity compared to the weak coupling as the $\lambda$ value increases. On the other hand, for comparison, the corresponding Schwarzschild solution (with $M=C_{0}$) and Kottler solution (with $M=C_{0},\Lambda=\Lambda_{\text{CDM}}$) are shown in the inset of the left panel. As noted and predicted in the previous section (for $\lambda>0$), the non-cosmological horizon is smaller than those of the Schwarzschild and Kottler solutions. In the right panel of Fig. 2, an illustrative solution with only one horizon (the cosmological horizon) is shown. Although the profile is regular at the origin, this solution contains a naked singularity, as the Ricci scalar diverges at $r=0$ (see the inset). Our numerical results indicate the existence of this type of solution when $\lambda\gtrapprox 0.09$ (with $C_{0},\Lambda_{\text{Eff}}$ fixed to the values specified in the figure). Finally, the approximate near horizon solution Eq. (20) is not shown because it is not valid in this branch (see Fig. 1).
Figure 3 presents a similar analysis (again fixing $\Lambda_{\text{eff}}$ to the cosmological constant value), but this time for a negative coupling constant $\lambda=-5$. The left panel shows an illustrative (non-physical) solution to demonstrate the methodology used: connecting the near and asymptotic regions by the numerical solution (black line). The region near the horizon is described by the solution Eq. (20) (blue line), while for large $r$, the solution matches the asymptotic approximation Eq. (15) (red line). Our results indicate (and validate our analytical findings) that, for this branch, the exterior horizons are larger than those of the respective Schwarzschild and Kottler solutions (see the corresponding insets). The right panel shows the results obtained by fixing the constant $C_{0}$ to the mass of SgrA.
@@
III Possible signatures of the cubic contribution in light deflection
Figure 4: Schematic and illustrative null geodesic triangular array. The left panel displays a schematic representation of a triangular array with vertices $\phi_{p}$, bounded by three null geodesics $\mathcal{C}_{p}$ ($p=1,2,3$) localized in the equatorial plane of a Riemannian space $\mathcal{M}_{\text{opt}}$ defined by the optical metric. Each geodesic $\mathcal{C}_{p}$ is associated with an impact parameter $b_{p}$ at an angular position $\phi_{b_{p}}$. For all triangular arrays explored in this paper, the vertices are fixed at: $\phi_{1}=\pi-\phi_{\text{Min}},\phi_{2}=\phi_{\text{Min}}$, and $\phi_{3}=\pi/2$, where $\phi_{\text{Min}}=0.35$rad. The corresponding internal angles $\beta_{p}$, are used to compute the angular difference Eq. (28). The central and right panels display triangle arrays where the spacetime under study is given by the numerical solutions of Eq. (II) with a positive coupling constant $\lambda=0.05$ (blue dashed line), and the background is described by the Kottler solution (solid black line). The central panel illustrates how the triangular array becomes more deformed due to its proximity to the gravitational source. In contrast, the right panel shows the same triangular array positioned much farther from the gravitational source. The insets reveal that, although these triangles appear similar, they exhibit minor discrepancies attributable to the $\lambda$ effect. Further details and discussions are provided in the main text.
In this section, we examine the angular difference $\alpha$, as proposed in Ref. [22], within the framework of Einsteinian cubic gravity. As mentioned in the introduction, this angular difference provides an alternative method for determining the deflection angle in static, spherically symmetric, and non-asymptotically flat spacetimes [41]. The method involves calculating the internal angles of a closed polygon (specifically, a triangle) formed by null geodesics in the spacetime of interest, which encode the effects of local curvature, and comparing them to a reference polygon (triangle) constructed in a background spacetime. A schematic representation of the construction of this triangle is shown in the left panel of Fig. 4.
As derived in Ref. [22], the angle $\alpha$ is determined from the angular difference between the internal angles of the triangular array, as follows:
@@ III Possible signatures of the cubic contribution in light deflection
where $\beta_{p}^{(i)}$ are the internal angles of the triangle placed in the spacetimes: $i=1$ the spacetime of interest and $i=2$ the spacetime that serves as background. For instance, if $i=1$ is the Schwarzschild spacetime and $i=2$ is Minkowski, then in the limit where the impact parameters of the null geodesics $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ (see Fig. 4) become sufficiently large (i.e., $b_{2}=b_{3}\to\infty$), the textbook formula for the deflection angle of the Schwarzschild metric is recovered.666 To review the procedure for obtaining the textbook formula for the deflection angle of the Schwarzschild metric, we refer the reader to Ref. [42], Section 7.17.17.1 of Ref. [43], or Section 555 of Ref. [44]. That is, the triangular configuration behaves as if the light source and the observer were located in the flat region at spatial infinity [22].
@@ III.1 Triangular numerical construction
where $b:=L/E=[r^{2}/f(r)]d\phi/dt$ is the impact parameter, defined in terms of the conserved photon energy $E=f(r)dt/d\lambda$ and angular momentum $L=r^{2}d\phi/d\lambda$, with $\lambda$ an affine parameter along the light trajectory. In terms of the dimensionless variables (27) and introducing the substitution $u:=1/r$, Eq. (29) can be rewritten as
@@
The second step consists in solving Eq. (30) with appropriate boundary conditions for the three null geodesics $C_{1},C_{2}$, and $C_{3}$, such that a triangular configuration is formed. For simplicity, the vertices are chosen as
@@
as shown in the left panel of Fig. 4. For the geodesic $\mathcal{C}_{1}$, we apply a Neumann boundary condition, $du/d\phi=0$ at $\phi=\pi/2$. For the remaining geodesics, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$, we use Dirichlet boundary conditions such that the bottom vertex of $\mathcal{C}_{2}$ ($\mathcal{C}_{3}$) coincides with the right (left) vertex of $\mathcal{C}_{1}$ at $\phi_{2}=\phi_{\text{Min}}$ ($\phi_{3}=\phi_{\text{Max}}$), where $[\phi_{\text{Min}},\phi_{\text{Max}}]$ is the angular domain of integration for $\mathcal{C}_{1}$ 888 The choice of these boundary conditions is meant to construct a simple approximation of the geodesics. Thus, the geodesic $C_{1}$ determines the boundary conditions for the other two, and its own boundary condition depends on the physical scenario being described (e.g. impact parameter, compact object).. For simplicity, the impact parameters $b_{2}$ and $b_{3}$ of geodesics $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ respectively, are taken to be equal and fixed to $1.5b_{1}$. We only vary the parameter $b_{1}$, which corresponds to the geodesic $\mathcal{C}_{1}$ that describes the closest trajectory to the gravitational source. The numerical integration is carried out using the same methodology as in Section II.6.
@@
Once the triangles are constructed, we compute their internal angles $\beta_{p}$ and, from these, the angular difference $\alpha$ as the final step. The internal angles are obtained from Ref. [22], as follows:
@@
III.2 Cubic terms contribution
@@
$b_{1}[r_{s}]$ 5.0-5.0- 5.0 1.5-1.5- 1.5 0.050.050.05 0.10.10.1
101010 $6.46\times 10^{3}$ $5.65\times 10^{3}$ $0.61\times 10^{2}$ $1.22\times 10^{2}$
@@
As initial examples, we considered as the spacetime of interest the (numerical and approximated) solutions of Einsteinian cubic gravity with the integration constant $C_{0}=4.3\times 10^{6}M_{\odot}$ (corresponding to the observed mass of SgrA) and coupling constant values: $\lambda=-5.0,-1.5,0.05,0.1$. The impact parameters of the triangular array were chosen such that the relation $b_{2}=b_{3}=1.5b_{1}$ holds, while $\phi_{\text{min}}$ was fixed at 0.30.30.3 rad. Table 2 presents the angular difference $\alpha$ computed using the Kottler (and Schwarzschild) solution as the background. As noted above, in this scenario, the values of $\alpha$ encode the (gravitational) effect of the cubic terms due to the mass contribution. Our results indicate that as $|\lambda|\to 0$ (for a fixed $b_{1}$), the value of $\alpha$ either diminishes or remains constant. This behavior highlights that the gravitational contribution of the cubic terms due to a massive source becomes negligible as $|\lambda|$ decreases. In the limit $\lambda=0$, it is expected that $\alpha=0$. Consequently, the Kottler/Schwarzschild solution is recovered, corresponding to the regime where general relativity is restored. The cases where $\alpha$ remains constant despite the value of $\lambda$ correspond to triangular arrays that are sufficiently far from the gravitational source (e.g. $b_{1}=10^{4}\,r_{s}$), such that the correction introduced by the coupling constant term in the asymptotic solution, Eq. (15), becomes negligible. In contrast, the value of the angular difference increases considerably when the coupling constant $\lambda$ is fixed and the impact parameter $b_{1}$ decreases. This is a direct consequence of the triangular configurations being closer to the gravitational source, which makes the contribution of the cubic terms more relevant.
@@
As a second example, we fixed the impact parameters $b_{1}=4r_{s}$, $b_{2}=b_{3}=1.5b_{1}$, and varied the coupling constant $\lambda$ within the interval [0,10][0,10][ 0 , 10 ] taking $C_{0}=4.3\times 10^{6}M_{\odot}$. The results are shown in Fig. 5. As observed from the left panel, when $\lambda\gtrapprox 0.1$, the solution does not present a black hole horizon (although the cosmological horizon remains present). This indicates that the value of $\lambda$ is constrained to $\lambda\lessapprox 0.1$ in order to ensure the presence of an exterior event horizon in black holes like SgrA (keeping in mind that the mass that we are using is very close to that of SgrA). The right panel shows the angular difference for these solutions. As observed, this increases as $\lambda$ grows. However, the results seem to indicate that it asymptotically approaches a certain value. We do not delve deeper into identifying it, as it corresponds to non-physically viable values of $\lambda$ due to the large modifications to the spacetime metric near the massive source, leading even to the disappearance of the event horizon.
@@
A similar analysis to the previous one is presented in Fig. 6, but in this case for $\lambda\in[-10,0]$ and $b_{1}=10r_{s}$. The last change is due to the exterior horizon moving to the right for this branch, and we need to ensure that the triangular array is not inside it. In this case, we always have an exterior event horizon (greater than $2C_{0}/M_{\textrm{Pl}}^{2}$), which could, in principle, reproduce black hole types like SgrA. As observed from the left panel, when $\lambda$ decreases (increases in magnitude), the horizon’s radius shifts to the right. However, this shift becomes progressively smaller; for example, the horizon displacement from a solution with $\lambda=-1$ to a solution with $\lambda=-2$ is approximately $0.2r_{s}$, while for $\lambda=-9$ to $\lambda=-10$ it is approximately $0.04r_{s}$. The right panel shows the angular difference. In this case, our results seem to indicate a linear increase (the $y$-axis is on a logarithmic scale). This is a direct consequence of the horizon approaching to the position of the triangular array. As discussed in the first example, this implies that the effect of the cubic terms becomes more relevant, and the differences relative to the triangular background array become more significant.